Circular No. 6518

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

γ-RAY BURSTS C. Meegan, V. Connaughton and G. Fishman, NASA Marshall Space

Flight Center; R. M. Kippen, University of Alabama at Huntsville; C. Kouveliotou, Universities Space Research Association; K. Hurley, University of California at Berkeley; T. Cline, D. Palmer, S. Barthelmy, P. Butterworth, B. Teegarden, H. Seifert and J. in’t Zand, NASA Goddard Space Flight Center; and E. Mazets and S. Golenetskij, Ioffe Institute of Physics and Technology, report: “Several spacecraft have detected a sequence of γ-ray bursts over a two-day period with locations that are consistent with a single source. The following positions (equinox 2000.0) have been determined by BATSE: Oct. 27.48897 UT, α = 4^{h}30^{m}, δ = −42^{o}.4, uncertainty 5^{o}.6, duration 100 s; 27.50141, 4^{h}35^{m}, −54^{o}.3, 5^{o}.8, 0.9 s (detection by Konus); 29.27404, 3^{h}58^{m}, −52^{o}.6, 4^{o}.6, 30 s (Konus); 29.28183, 3^{h}59^{m}, −48^{o}.9, 0^{o}.3, 750 s (Ulysses, TGRS and Konus). The proximity (i.e., the quadratic sum of the statistical errors, excluding BATSE’s systematic error of ∼ 1^{o}.6) of the first three BATSE event locations to the fourth, including Interplanetary (IPN) locations using BATSE, Ulysses and Konus, are 1^{o}.4, 3^{o}.0 and 1^{o}.0, respectively. The Ulysses-BATSE IPN annulus is described by a center at α = 23^{h}46^{m}.63, δ = −31^{o}24′.5 with a radius of 49^{o}.883, fullwidth 0^{o}.078. The much wider Ulysses-Konus annulus defines the end points of the Ulysses-BATSE annular segment at α = 3^{h}58^{m}.9, δ = −47^{o}04′ and α = 4^{h}01^{m}.3, δ = −52^{o}35′ with a 99.7-percent confidence level. A soft γ repeater can be excluded as a common source by the durations of the events, as well as by their spectra, which are consistent with classical γ-ray bursts. The temporal structures of the events are quite different, apparently ruling out a gravitational-lens interpretation. The third and fourth events are very probably separate triggers from a single burst, making the combined event the longest burst (1420 seconds) ever seen in this energy range. Although a posteriori calculations are problematic, the probability of such a temporal and spatial coincidence of four unrelated events is low.”

COMET 22P/KOPFF Total visual magnitudes (L = reflector): Aug. 11.96 UT, 8.8 (J. Carva-

jal, Teruel, Spain, 0.21-m L); 22.23, 9.5 (C. E. Spratt, Victoria, BC, 0.20-m L); Sept. 3.89, 8.9 (J. M. San Juan, Madrid, Spain, 0.21-m L); 17.87, 9.6 (J. M. Trigo, Castellon, Spain, 0.17-m L); Oct. 3.11, 11.2 (A. Hale, Cloudcroft, NM, 0.41-m L); 9.17, 12.1 (R. Keen, Mt. Thorodin, CO, 0.32-m L); Nov. 4.75, 13.0 (M. Plsek, Lelekovice, Czech Republic, 0.35-m L).

1996 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden